carmen is twice tom's age.3 years ago carmen was 3 times as old as tom.how old is carmen?how old is tom?
Answer by stanbon(75887)

carmen is twice tom's age.3 years ago carmen was 3 times as old as tom.how old is carmen?how old is tom?
----
Equations:
c = 2t
(c-3) = 3(t-3)
---
Substitute for "c" and solve for "t":
(2t-3) = 3t-9
t = 6 (tom's age now)
---
Solve for "c":
c = 2t
c = 12 (camen's age now)
===========================
Cheers,
Stan H.